hayhay6:

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = −6 3x − 4y = −18 Options : 6x − y = −15 3y = 9 6x − 8y = −36 −6y = −42 8x + 4y = −4 14x = −10 12x + 4y = −12 15x = −30

Vocaloid:

You’ll need to solve the first system, determine the solution, then keep solving the other ones until you find one that has the same solution as the first one For 6x + 2y = −6 3x − 4y = −18 You can multiply the first equation by 2, ending up with 12x + 4y = -12, and from there you can add the two equations to eliminate y, then solve for x, then plug x back into one of the original equations to solve for y.